# Enhancing Answer Reliability Through Inter-Model Consensus of Large Language Models

### v1

@@ Abstract
We explore the collaborative dynamics of an innovative language model interaction system involving advanced models such as GPT-4-0125-preview, Meta-LLaMA-3-70B-Instruct, Claude-3-Opus, and Gemini-1.5-Flash. These models generate and answer complex, PhD-level statistical questions without exact ground-truth answers. Our study investigates how inter-model consensus enhances the reliability and precision of responses. By employing statistical methods such as chi-square tests, Fleiss’ Kappa, and confidence interval analysis, we evaluate consensus rates and inter-rater agreement to quantify the reliability of collaborative outputs. Key results reveal that Claude and GPT-4 exhibit the highest reliability and consistency, as evidenced by their narrower confidence intervals and higher alignment with question-generating models. Conversely, Gemini and LLaMA show more significant variability in their consensus rates, as reflected in wider confidence intervals and lower reliability percentages. These findings demonstrate that collaborative interactions among large language models (LLMs) significantly improve response reliability, offering novel insights into autonomous, cooperative reasoning and validation in AI systems.
Keywords: Large Language Models, Collaborative Intelligence, Answer Validation, Game Theory, Statistical Analysis
@@ 1 Introduction
The emergence of large language models (LLMs) has transformed natural language processing and artificial intelligence, enabling machines to perform complex language tasks with unprecedented proficiency. Models such as OpenAI’s GPT-4, Meta’s LLaMA series, Anthropic’s Claude, and Google’s Gemini have demonstrated remarkable capabilities in generating human-like text, understanding context, and even exhibiting reasoning abilities Mann et al. (2020); Touvron et al. (2023). These advancements have opened new avenues for automated knowledge generation and validation in specialized domains Kojima et al. (2022).
However, a significant challenge remains in validating the outputs of these models when predefined correct answers or ground truths are unavailable. This issue is particularly critical in specialized fields like advanced statistics, where the complexity and nuance of questions cause manual verification to be resource-intensive and impractical Jiao et al. (2023). Traditional validation methods that rely on comparing model outputs against predetermined correct answers become insufficient in such contexts Hendrycks et al. (2020).
1.1 Research Context and Motivation
Leveraging multiple LLMs to enhance answer reliability through collaboration and consensus presents a promising approach. By harnessing the collective intelligence of several models, it becomes possible to approximate correctness and identify consensus even in the absence of ground truth He et al. (2022). This methodology draws inspiration from ensemble learning techniques in machine learning Dietterich (2000) and the wisdom of crowds phenomenon in human collective intelligence Mennis (2006).
@@ 1.2 Theoretical Framework
The Distributed Cognition Framework Hutchins (1995) offers theoretical grounding for understanding how cognitive processes can be distributed across multiple artificial agents. This framework is particularly relevant because different models may encode complementary aspects of statistical knowledge, leading to more comprehensive problem-solving capabilities Du et al. (2024). Complex statistical reasoning tasks can be decomposed and processed across multiple models, potentially improving overall solution quality Liu et al. (2024). Furthermore, the framework helps explain how information flows between models during collaborative problem-solving Naik (2024).
Mathematical models of consensus formation provide the theoretical basis for understanding how agreement emerges among multiple decision-making agents Friedkin (1990). These models inform the fundamental dynamics of collaborative AI systems through three key aspects: the dynamics of agreement, which describe how consensus evolves; weighted influence mechanisms, which define the relative impact of each agent in the consensus process Olfati-Saber et al. (2007); and the conditions for convergence that ensure the stability of collaborative solutions Baronchelli (2018). Recent work by Bahrami et al. (2010) has demonstrated how these theoretical frameworks can be effectively applied to multi-agent decision-making systems, offering a robust foundation for understanding collaborative behavior in artificial intelligence.
@@
The validation methodology emerging from this integrated framework emphasizes the importance of multiple perspectives in evaluating model outputs Kittur and Kraut (2008). By considering various theoretical viewpoints, we can better assess the reliability and robustness of collaborative AI solutions Malone and Bernstein (2022). This comprehensive theoretical foundation grounds our investigation in established principles and provides new insights into the unique challenges posed by AI collaboration in statistical reasoning tasks Page (2008).
1.3 Research Gap
While existing research has extensively examined individual LLM performance Ahn et al. (2024) and basic ensemble methods Huang et al. (2024), there is a notable gap in understanding how multiple state-of-the-art LLMs can collaboratively validate complex knowledge in the absence of ground truth. This gap is particularly significant in specialized academic domains where:
  • The complexity of questions requires sophisticated reasoning capabilities.
  • Manual validation by human experts is time-consuming and costly.
  • Traditional automated validation methods are insufficient.
  • The dynamic nature of knowledge makes maintaining up-to-date ground truths challenging.
Addressing these challenges necessitates novel approaches that leverage collaborative validation among LLMs, providing a pathway toward reliable, scalable, and efficient knowledge validation.
1.4 Objectives of the Research
This study aims to explore and understand the collaborative dynamics among different LLMs in statistical reasoning tasks. The primary objectives are to examine how answers from different LLMs align or differ when responding to complex statistical questions generated by one of the models. This includes investigating how different LLMs complement each other in collaborative statistical reasoning and understanding the patterns of their interactions. Then, develop a framework for evaluating the reliability of answers based on model consensus, especially in the absence of ground truth. This involves quantifying the reliability of collaborative outcomes without predefined correct answers and identifying metrics that best capture the quality of inter-model agreement. Also, statistical techniques such as chi-square tests and Fleiss’ Kappa are used to assess the significance of consensus rates and measure inter-model agreement. This objective supports validating collaborative mechanisms and the effectiveness of consensus formation among multiple LLMs.
1.5 Significance and Contributions
This research makes several significant contributions to the field:
  1. 1.
    Introduces a novel framework for validating complex knowledge through multi-model collaboration.
  2. 2.
    Provides empirical evidence for the effectiveness of collaborative validation approaches.
  3. 3.
    Establishes benchmarks for future research in LLM-based knowledge validation.
The findings of this study have implications in several areas, such as educational technology and automated assessment systems, research validation in specialized academic domains, development of more reliable AI-powered knowledge systems, and improving our understanding of how artificial systems can collaborate and exhibit collective intelligence to solve complex problems to name a few.
@@ 2 Literature Review
AI trends have moved from advances in image processing and computer vision Voulodimos et al. (2018); Azad et al. (2024) to progress in reinforcement learning Wiering and Van Otterlo (2012); Jebellat et al. (2021, 2024) and natural language processing Chowdhary and Chowdhary (2020); Pillai (2023), leading to today’s focus on LLMs and generative AI.
LLMs have achieved significant milestones in natural language understanding and generation Radford et al. (2019). The GPT series, notably, has demonstrated remarkable capabilities in few-shot and zero-shot learning scenarios Achiam et al. (2023). Recent advancements in model architectures and training approaches have led to increasingly sophisticated systems, such as Claude-3 Anthropic (2024) and Gemini Team et al. (2023), which exhibit enhanced reasoning abilities in specialized domains.
The concept of collaborative intelligence among AI agents has emerged as a promising approach for enhancing problem-solving capabilities Dafoe et al. (2020). Recent studies have shown that model collaboration can significantly improve reasoning capabilities through cross-model validation Yin et al. (2023). Additionally, it enhances robustness, particularly in tackling complex problem-solving tasks Sun et al. (2023); Jebellat and Jebellat (2024); Gholami Davoodi et al. (2024), and leads to more reliable outputs by leveraging consensus mechanisms.
Integrating collaborative strategies in artificial intelligence (AI) systems has consistently demonstrated enhanced performance across various domains. For instance, Lu et al. (2024) conducted a comprehensive survey highlighting that merging, ensembling, and cooperative approaches among LLMs lead to superior outcomes in natural language processing tasks. In the medical field, Rezk and Selim (2024) reviewed metaheuristic-based ensemble learning methods, emphasizing their effectiveness in improving diagnostic accuracy and treatment planning. Similarly, in mathematical problem-solving, researchers at MIT developed a multi-AI collaboration framework that enhances reasoning and factual accuracy in LLMs, resulting in more reliable solutions to complex mathematical queries MIT News (2023). In software engineering, ensemble methods have been shown to improve code defect detection and software quality assurance processes Gupta et al. (2022). These findings collectively highlight the significant advantages of collaborative AI approaches, particularly as LLMs continue to evolve and scale.
@@ 3 Methodology
3.1 Research Design
This study employs a mixed-methods approach to investigate the collaborative dynamics among LLMs in statistical reasoning tasks. The experimental design incorporates quantitative analysis of model consensus patterns. The study framework can be formalized as follows:
@@ 3.1 Research Design
3.2 Overview
We examine how multiple LLMs collaborate to generate and validate complex, PhD-level statistical multiple-choice questions (MCQs) without ground-truth answers. A total of $N=100$ MCQs were generated and answered. One LLM acted as the question generator for each question, while the remaining three independently provided answers and justifications. Question generator and answerer roles were rotated among the models to ensure interchangeability and mitigate model-specific biases.
3.3 Language Models Used
We utilized four state-of-the-art LLMs, each with distinct architectural characteristics:
@@ 3.3 Language Models Used
Each model was accessed via its respective API, using default hyperparameters unless specified otherwise. All models operated under identical conditions to maintain consistency.
3.4 Question Generation and Answering Processes
To ensure the creation of challenging and diverse multiple-choice questions (MCQs) suitable for PhD-level statistics, we designed an integrated framework combining question generation and independent answering LLMs. This framework also incorporated strategies to mitigate potential biases in the question generation and answering processes.
The question generation phase leveraged a comprehensive concept map of advanced statistical topics, as illustrated in table 2. This table encompasses a diverse set of topics $\mathcal{T}$ and subtopics $\mathcal{S}$. For each question, a topic $t_{i}\in\mathcal{T}$ and subtopic $s_{i}\in\mathcal{S}$ were randomly selected to ensure uniform coverage and diversity. The question-generating LLM received a carefully designed prompt $P_{q}$:
”Generate a challenging PhD-level multiple-choice question in the field of [Topic], focusing on [Specific Concept]. The question should have four answer options labeled A, B, C, and D, with only one correct answer. Ensure the question tests deep understanding and critical thinking skills.”
Here, the placeholders [Topic] and [Specific Concept] were populated based on the selected topic $t_{i}$ and subtopic $s_{i}$ from the table 2. The generated output included the question $Q_{i}$, the answer set $\{A_{i},B_{i},C_{i},D_{i}\}$, the correct answer $A_{i}^{c}$, and an explanation $E_{i}$. Notably, $A_{i}^{c}$ and $E_{i}$ were retained for analysis but withheld from the answering models to replicate real-world conditions where ground-truth answers are unavailable.
To mitigate potential biases, we implemented several strategies. First, topic diversity was ensured by including various topics and subtopics, preventing overrepresenting specific areas. Second, the crafted prompts were neutral and avoided introducing leading language or biases, ensuring no influence on the models’ responses in a particular direction. Finally, all generated content was manually reviewed to identify and exclude inappropriate or biased material, thereby preserving the integrity and fairness of the dataset.
@@ 3.4 Question Generation and Answering Processes
Time series analysis ARIMA models, Stationarity, Seasonality, Spectral analysis
@@
In the answering phase, three distinct LLMs independently attempted the generated questions. Each answering model received the question $Q_{i}$ and answer set $\{A_{i},B_{i},C_{i},D_{i}\}$ alongside the following prompt $P_{a}$:
@@
Each model independently selected an answer $a_{ij}$ for question $Q_{i}$, where $j\in\{1,2,3\}$ indexes the answering models, and provided a justification $J_{ij}$. The models were isolated from one another, ensuring no exchange of information about answers or justifications to prevent bias or collusion. This isolation, combined with the neutral prompts and diverse question topics, further minimized bias in the answering process.
3.5 Inter-Model Consistency Analysis
@@
This dataset encompassed the generated questions $Q_{i}$, their corresponding multiple-choice options $\{A_{i},B_{i},C_{i},D_{i}\}$, the responses from each participating LLM—including selected answers $a_{ij}$ and justifications $J_{ij}$—and metadata such as model identifiers, timestamps, and relevant model-specific parameters. This structured approach facilitated quantitative and qualitative analysis of the reasoning processes and agreement patterns among models Bommasani et al. (2021).
We analyzed the alignment and divergence in the selected answers $a_{ij}$ to quantify inter-model consistency. The degree of consensus was categorized as follows:
  • Full Agreement: All three models selected the same answer.
  • Partial Agreement: Two models selected the same answer, while one differed.
  • No Agreement: All models selected different answers.
This classification provided a robust framework for evaluating agreement levels and understanding the variability in model decision-making, offering insights into their collective and individual reasoning patterns.
3.6 Answer Validation Mechanisms
@@
To determine the consensus answer $A_{i}^{\text{cons}}$ for a given question $Q_{i}$, we utilized a majority voting mechanism over the responses from multiple LLMs. Formally, the consensus answer is defined as:
@@
This formulation assigns a frequency-based score to each answer option $k$, with the majority-selected response designated as the consensus answer. Both partial and complete agreements are considered valid, forming the basis for subsequent analyses.
In the absence of ground truth labels, we introduced a reliability metric to assess the trustworthiness of the consensus answers. Reliability measures the alignment between the consensus response and the answer provided by the LLM that generated the question. Let $A_{i}^{\text{LLM-q}}$ denote the querying LLM’s response for question $Q_{i}$. The reliability score $R_{i}$ is computed as:
@@
A higher reliability score indicates a more substantial alignment between the consensus answer and the querying LLM’s response, increasing confidence in the consensus’s validity. This measure is precious in settings where ground truth data is unavailable.
To evaluate the robustness of consensus rates across LLMs, we calculated confidence intervals (CIs) using a bootstrap resampling approach. Confidence intervals provide a probabilistic range for the actual consensus rate, capturing variability in the observed data. Narrow CIs indicate high reliability, while wider CIs reflect more significant uncertainty.
Bootstrap resampling enables CI estimation without assumptions about the underlying data distribution. Specifically, $B$ bootstrap samples are generated by sampling the original dataset with replacement. The mean agreement rate is computed for each sample, constructing a bootstrap distribution of the consensus rate. The $\alpha/2$-th and $21-\alpha/$ 2-th percentiles of this distribution define the lower and upper CI bounds, respectively. For a 95% CI, these correspond to the 2.5th and 97.5th percentiles.
By incorporating confidence intervals, we mitigate uncertainty in agreement rates and establish a statistical basis for comparing LLMs. Overlapping CIs suggest no significant differences between models, whereas non-overlapping intervals indicate statistically significant differences. This methodology ensures a robust and interpretable evaluation of model reliability, mainly when ground truth labels are unavailable.
To further validate reliability and assess statistical significance, we employed the following additional statistical measures:
@@
A chi-square test was used to evaluate whether the distribution of selected answers deviated significantly from random chance. The test statistic was calculated as:
@@
  • $O_{k}$ is the observed frequency of answer choice $k$ across all models and questions.
  • $E_{k}$ is the expected frequency of answer choice $k$, assuming uniform random selection, $E_{k}=\frac{N\times n_{j}}{K}$.
  • $K=4$ is the number of answer choices.
@@
  • $n_{j}=3$ is the number of answering models.
@@
The value of $\kappa$ ranges from $-1$ (perfect disagreement) to $1$ (perfect agreement), with 00 indicating no agreement beyond chance.
These mechanisms collectively ensured a robust framework for validating model responses and assessing the reliability of the generated consensus.

### v2

@@ Abstract
We propose a collaborative framework in which multiple large language models—including GPT-4-0125-preview, Meta-LLaMA-3-70B-Instruct, Claude-3-Opus, and Gemini-1.5-Flash—generate and answer complex, PhD-level statistical questions when definitive ground truth is unavailable. Our study examines how inter-model consensus improves both response reliability and identifies the quality of the generated questions. Employing chi-square tests, Fleiss’ Kappa, and confidence interval analysis, we quantify consensus rates and inter-rater agreement to assess both response precision and question quality. Key results indicate that Claude and GPT-4 produce well-structured, less ambiguous questions with a higher inter-rater agreement, as shown by narrower confidence intervals and greater alignment with question-generating models. In contrast, Gemini and LLaMA exhibit greater variability and lower reliability in question formulation. These findings demonstrate that collaborative interactions among large language models enhance response reliability and provide valuable insights for optimizing AI-driven collaborative reasoning systems.
Keywords: Large Language Models, Collaborative Intelligence, Inter-Model Consensus, Statistical Analysis, Confidence Interval
@@ 1 Introduction
The emergence of large language models (LLMs) has transformed natural language processing and artificial intelligence (AI), enabling machines to perform complex language tasks with unprecedented proficiency. Models such as OpenAI’s GPT-4, Meta’s LLaMA series, Anthropic’s Claude, Google’s Gemini, and DeepSeek have demonstrated remarkable capabilities in generating human-like text, understanding context, and even exhibiting reasoning abilities Mann et al. (2020); Touvron et al. (2023); Guo et al. (2025). These advancements have opened new avenues for automated knowledge generation and validation in specialized domains Kojima et al. (2022).
However, two significant challenges remain: validating model outputs when predefined correct answers or ground truths are unavailable, and assessing the quality of the questions these models generate. These issues are especially critical in specialized fields like advanced statistics, where the complexity of questions makes manual verification resource-intensive and impractical Jiao et al. (2023). In such contexts, traditional validation methods that compare outputs against predetermined answers are simply insufficient Hendrycks et al. (2020).
Leveraging multiple LLMs to enhance the reliability of the answer and assess the quality of the questions through collaboration and consensus presents a promising approach. By utilizing the collective intelligence of several models, it becomes possible to approximate correctness and identify consensus even in the absence of ground truth He et al. (2022). Additionally, this strategy provides insights into question quality by evaluating how consistently other models respond. This methodology draws inspiration from ensemble learning techniques in machine learning Dietterich (2000) and the wisdom of crowds phenomenon in human collective intelligence Mennis (2006).
@@
The Distributed Cognition Framework Hutchins (1995) offers theoretical grounding for understanding how cognitive processes can be distributed across multiple artificial agents. This framework is particularly relevant because different models may encode complementary aspects of statistical knowledge, leading to more comprehensive problem-solving capabilities Du et al. (2024). Complex statistical reasoning tasks can be decomposed and processed in multiple models, potentially improving the overall quality of the solution Liu et al. (2024); Naik (2024).
Mathematical models of consensus formation provide the theoretical basis for understanding how agreement emerges among multiple decision-making agents Friedkin (1990). These models inform the fundamental dynamics of collaborative AI systems through three key aspects: the dynamics of agreement, which describe how consensus evolves; weighted influence mechanisms, which define the relative impact of each agent in the consensus process Olfati-Saber et al. (2007); and the conditions for convergence that ensure the stability of collaborative solutions Baronchelli (2018). Work by Bahrami et al. (2010) has demonstrated how these theoretical frameworks can be effectively applied to multi-agent decision-making systems, offering a robust foundation for understanding collaborative behavior in AI.
@@
The validation methodology emerging from this integrated framework emphasizes the importance of multiple perspectives in evaluating model outputs Kittur and Kraut (2008). By considering various theoretical viewpoints, we can better assess the reliability and robustness of collaborative AI solutions Malone and Bernstein (2022). This comprehensive theoretical foundation supports our investigation of established principles and provides new insights into the unique challenges posed by AI collaboration in statistical reasoning tasks Page (2008).
Although existing research has extensively examined individual LLM performance Ahn et al. (2024) and basic ensemble methods Huang et al. (2024), there is a notable gap in understanding how multiple state-of-the-art LLMs can collaboratively validate complex knowledge in the absence of ground truth. This gap is particularly significant in specialized academic domains where the complexity of questions demands sophisticated reasoning capabilities, manual validation of both question quality and responses by human experts is time-consuming and costly, and traditional automated validation methods are insufficient. Furthermore, the dynamic nature of knowledge makes maintaining up-to-date ground truths a formidable challenge. Addressing these issues requires leveraging collaborative validation among LLMs, which provides a pathway toward reliable, scalable, and efficient knowledge validation.
This study aims to explore and understand the collaborative dynamics among different LLMs in statistical reasoning tasks. Our primary objective is to examine how answers from various LLMs align or differ when responding to complex statistical questions generated by one of the models, thereby revealing how these models complement each other and offering insights into the quality of their question-generation capabilities. We further develop a framework for evaluating answer reliability based on model consensus in the absence of a predefined ground truth. To achieve this, we quantify the reliability of collaborative outcomes and identify metrics that best capture inter-model agreement, employing statistical techniques such as chi-square tests and Fleiss’ Kappa to assess the significance of consensus rates. These efforts support validating collaborative mechanisms and the effectiveness of consensus formation among multiple LLMs.
The findings of this study have the potential to influence educational technology, automated assessment, research validation in specialized fields, and the development of more robust AI-powered knowledge systems while enhancing our understanding of collaborative AI.
@@ 2 Literature Review
AI trends have moved from advances in image processing and computer vision Voulodimos et al. (2018); Azad et al. (2024) to progress in reinforcement learning Wiering and Van Otterlo (2012); Jebellat et al. (2024) and natural language processing Chowdhary and Chowdhary (2020); Pillai (2023), leading to today’s focus on LLMs and generative AI.
LLMs have achieved significant milestones in natural language understanding and generation Radford et al. (2019). The GPT series, notably, has demonstrated remarkable capabilities in few-shot and zero-shot learning scenarios Achiam et al. (2023). Recent advancements in model architectures and training approaches have led to increasingly sophisticated systems, such as Claude-3 Anthropic (2024), Gemini Team et al. (2023) and DeepSeek Guo et al. (2025), which exhibit enhanced reasoning abilities in specialized domains.
The concept of collaborative intelligence among AI agents has emerged as a promising approach for enhancing problem-solving capabilities Dafoe et al. (2020). Recent studies have shown that model collaboration can significantly improve reasoning capabilities through cross-model validation Yin et al. (2023). Additionally, it enhances robustness, particularly in tackling complex problem-solving tasks Sun et al. (2023); Davoodi et al. (2024), and leads to more reliable outputs by leveraging consensus mechanisms.
Integrating collaborative strategies in AI systems has consistently demonstrated performance enhancement across various domains. For instance, Lu et al. (2024) conducted a comprehensive survey highlighting that merging, ensembling, and cooperative approaches among LLMs lead to superior outcomes in natural language processing tasks. In the medical field, Rezk and Selim (2024) reviewed metaheuristic-based ensemble learning methods, emphasizing their effectiveness in improving diagnostic accuracy and treatment planning. Similarly, in mathematical problem-solving, researchers at MIT developed a multi-AI collaboration framework that enhances reasoning and factual accuracy in LLMs, resulting in more reliable solutions to complex mathematical queries MIT News (2023). In software engineering, ensemble methods have been shown to improve code defect detection and software quality assurance processes Gupta et al. (2022). These findings collectively highlight the significant advantages of collaborative AI approaches, particularly as LLMs continue to evolve and scale.
@@ 3 Methodology
This study employs a mixed-methods approach to investigate the collaborative dynamics among LLMs in statistical reasoning tasks. The experimental design incorporates a quantitative analysis of model consensus patterns. The study framework can be formalized as follows:
@@
We examine how multiple LLMs collaborate to generate and validate complex, PhD-level statistical multiple-choice questions (MCQs) without ground-truth answers, while also assessing each LLM’s ability to produce meaningful questions. A total of $N=100$ MCQs were generated and answered. An LLM acted as the question generator for each question, while the remaining three independently provided answers and justifications. Question generator and answerer roles were rotated among the models to ensure interchangeability and mitigate model-specific biases.
We utilized four state-of-the-art LLMs, each with distinct architectural characteristics.
@@
Each model was accessed via its respective API, with standardized key parameters to ensure comparable responses. All models operated under identical conditions to maintain consistency.
To ensure the creation of challenging and diverse multiple choice questions (MCQs) suitable for PhD-level statistics, we designed an integrated framework that combined question generation and independent answering LLMs. This framework also incorporated strategies to mitigate potential biases in the question generation and answering processes.
The question generation phase leveraged a comprehensive concept map of advanced statistical topics, as illustrated in table 2. This table covers a diverse set of topics $\mathcal{T}$ and subtopics $\mathcal{S}$. For each question, a topic $t_{i}\in\mathcal{T}$ and a subtopic $s_{i}\in\mathcal{S}$ were randomly selected to ensure uniform coverage and diversity. The LLM which generated questions received a carefully designed prompt $P_{q}$:
”Generate a challenging multi-choice question at the PhD level in the field of [Topic], focusing on [Specific Concept]. The question should have four answer options labeled A, B, C, and D, with only one correct answer. Ensure that the question tests deep understanding and critical thinking skills.”
Here, the placeholders [Topic] and [Specific Concept] were populated based on the selected topic $t_{i}$ and subtopic $s_{i}$ from the table 2. The generated output included the question $Q_{i}$, the set of answers $\{A_{i},B_{i},C_{i},D_{i}\}$, the correct answer $A_{i}^{c}$, and an explanation $E_{i}$. In particular, $A_{i}^{c}$ and $E_{i}$ were retained for analysis but withheld from the answer models to replicate real-world conditions where ground truth answers are not available.
To mitigate potential biases, we implemented several strategies. First, topic diversity was ensured by including various topics and subtopics, preventing over-representing specific areas. Second, the crafted prompts were neutral and avoided introducing leading language or biases, ensuring no influence on the models’ responses in a particular direction. Finally, all generated content was manually reviewed to identify and exclude inappropriate or biased material, thus preserving the integrity and fairness of the dataset.
@@
Time series analysis ARIMA models, Stationary, Seasonality, Spectral analysis
@@
In the answer phase, three distinct LLMs independently attempted the generated questions. Each answering model received the question $Q_{i}$ and answer set $\{A_{i},B_{i},C_{i},D_{i}\}$ alongside the following prompt $P_{a}$:
@@
Each model independently selected an answer $a_{ij}$ for the question $Q_{i}$, where $j\in\{1,2,3\}$ indexes the response models and provided a justification $J_{ij}$. The models were isolated from each other, ensuring that no information was exchanged about answers or justifications to prevent bias or collusion. This isolation, combined with the neutral prompts and diverse question topics, further minimized bias in the answering process.
@@
This data set encompassed the generated questions $Q_{i}$, their corresponding multiple choice options $\{A_{i},B_{i},C_{i},D_{i}\}$, the responses of each participating LLM - including selected answers $a_{ij}$ and justifications $J_{ij}$ and metadata such as model identifiers, timestamps, and relevant model-specific parameters. This structured approach facilitated the quantitative and qualitative analysis of reasoning processes and agreement patterns among models Bommasani et al. (2021).
We analyze the alignment and divergence in the selected answers $a_{ij}$ to quantify the consistency between models. The degree of consensus was categorized into three levels: full agreement, where all three models selected the same answer; partial agreement, where two models selected the same answer while one differed; and no agreement, where all models selected different answers. This classification provided a robust framework for evaluating agreement levels and understanding the variability in model decision-making, providing insights into their collective and individual reasoning patterns.
@@
To determine the consensus answer $A_{i}^{\text{cons}}$ for a given question $Q_{i}$, we used a majority voting mechanism over the responses of multiple LLMs. Formally, the consensus answer is defined as:
@@
This formulation assigns a frequency-based score to each answer option $k$, with the majority-selected response designated as the consensus answer. Both partial and complete agreements are considered valid and form the basis for subsequent analyses.
In the absence of ground truth labels, we introduced a reliability metric to assess the trustworthiness of the consensus answers. Reliability measures the alignment between the consensus response and the answer provided by the LLM that generated the question. Let $A_{i}^{\text{LLM-q}}$ denote the querying LLM’s response to question $Q_{i}$. The reliability score $R_{i}$ is computed as:
@@
A higher reliability score indicates a more substantial alignment between the consensus answer and the querying LLM’s response, increasing confidence in the consensus’s validity. This measure is particularly valuable in settings where ground truth data is unavailable, and, when combined with consensus analysis, it provides qualitative insights into the performance of models in question generation tasks.
To evaluate the robustness of the consensus rates in LLMs, we calculated confidence intervals (CI) using a bootstrap resampling approach. Confidence intervals provide a probabilistic range for the actual consensus rate, capturing variability in the observed data. Narrow CIs indicate high reliability, while wider CIs reflect greater uncertainty.
Bootstrap resampling enables CI estimation without assumptions about the underlying data distribution. Specifically, $B=1000$ bootstrap samples are generated by sampling the original data set with replacement. The mean agreement rate is computed for each sample, constructing a bootstrap distribution of the consensus rate. The $\alpha/2$-th and $21-\alpha/$ 2-th percentiles of this distribution define the lower and upper CI bounds, respectively. For a 95% CI, these correspond to the 2.5 and 97.5 percentiles.
By incorporating confidence intervals, we mitigate the uncertainty in agreement rates and establish a statistical basis for comparing LLMs. Overlapping CIs suggest that there are no significant differences between models, whereas nonoverlapping intervals indicate statistically significant differences. This methodology ensures a robust and interpretable evaluation of model reliability, mainly when ground truth labels are unavailable.
To further validate reliability and assess statistical significance, we employed the following additional statistical measures.
@@
A chi-square test was used to evaluate whether the distribution of selected responses deviated significantly from random chance. The test statistic was calculated as:
@@
  • $O_{k}$ is the observed frequency of the answer choice $k$ in all models and questions.
  • $E_{k}$ is the expected frequency of the answer choice $k$, assuming uniform random selection $E_{k}=\frac{N\times n_{j}}{K}$.
  • $K=4$ is the number of choice of responses.
@@
  • $n_{j}=3$ is the number of response models.
@@
The value of $\kappa$ ranges from $-1$ (perfect disagreement) to $1$ (perfect agreement), with 00 indicating that there is no agreement beyond chance.
These mechanisms collectively ensured a robust framework for validating model responses and assessing the reliability of the consensus generated.